OpenStudy (evies163):

Help! Please! Solve for x 4x+8=x+5+7x-4

OpenStudy (jackthegreatest):

4x+8=x+5+7x-4 4x - x - 7x = 5-4-8 -4x = -7 x = 7/4

OpenStudy (jackthegreatest):

@EvieS163

OpenStudy (shadowlegendx):

\[4x+8=x+5+7x-4 \] Like terms \[4x + 8 = 8x + 1\] Isolate x by getting the variable on one side, alone Subtract 8x from both sides \[-4x + 8 = 1\] Subtract 8 from both sides to isolate x \[-4x = -7\] Divide the entire equation by -4 \[x = \frac{ -4 }{ -7 } \rightarrow x = \frac{ 4 }{ 7 }\]

OpenStudy (shadowlegendx):

Remember, a negative divided by a negative, is a positive

OpenStudy (shadowlegendx):

Let me correct that last equation it should be, \[x = \frac{ -7 }{ -4 } \rightarrow x = \frac{ 7 }{ 4 }\]

OpenStudy (shadowlegendx):

My apologies

OpenStudy (mathmale):

Evie? Where are you? Please become involved in this work.